OpenStudy (anonymous):

Solve each equation or formula for the variable specified. -3x + b = 6x, for x You'll help if you're cool and/or have a large peni.s.

OpenStudy (uri):

@abb0t wants to halp in dat case...cuz i don't have one..:(

OpenStudy (abb0t):

add 3x to both sides, then divide by 9 to both sides.

OpenStudy (anonymous):

So the answer is 1?

OpenStudy (anonymous):

Or is it 9? Plz help oh wise one!

OpenStudy (anonymous):

-3x+b=6x if you add 3x to both sides then it becomes b=9x then you need to divide both sides by nine making it (1/9)b=x

OpenStudy (anonymous):

So it's 1/9b = x?

OpenStudy (anonymous):

So x is 9?

OpenStudy (anonymous):

not the second one you said the first one you said after I posted is right